# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v3

@@ 2.2 Pre-trained Foundation Model for Gaia XP Spectra
Gaia XP low-resolution spectra (De Angeli et al., 2023), obtained via the Blue (with resolving power $R$ ranging from 30 to 100) and Red ($R$ ranging from 70 to 100) Photometers (BP and RP), are essential for determining key stellar properties such as stellar atmospheric parameters and chemical compositions. The differences in resolution and wavelength range compared to LAMOST LRS motivate the construction of a separate foundation model tailored to the XP modality.
We pre-train this model using one million Gaia XP spectra, with around 80% having matching LAMOST LRS spectra. Each XP spectrum consists of 343 flux points spanning wavelengths from 336 nm to 1021 nm. For such low-resolution spectra, broad-band color information may sometimes carry more discriminative power than individual spectral features. We therefore normalize each spectrum by its flux at 550 nm. This wavelength lies near the center of the $V$-band, providing a stable reference point across diverse stellar types.
@@
Two types of models are explored for Gaia XP. The first is a transformer-based MT model, structurally similar to that used for LAMOST LRS, but with tokenization at the individual flux-point level, yielding 343 tokens per spectrum. Like the LRS model, this version uses 6 self-attention layers and masks 6 chunks (20 tokens each) during training.
@@
The second model is a simple MLP-based autoencoder (denoted as OAE), with a bottleneck layer of 768 dimensions. Both XP models have approximately the same number of trainable parameters as the LRS model and are trained for the same number of epochs. In this paper, we use the OAE for Gaia XP to test results in tables, unless otherwise noted, because of its better performance, as discussed in Section 5.6.
2.3 Contrastive Learning with Decoders
@@
To align the two spectral modalities, we use 820,568 paired LAMOST LRS and Gaia XP spectra. The backbone of the alignment model leverages contrastive loss, augmented with auxiliary decoders that contribute additional supervision.
@@
As shown in Figure 1, embeddings from the LRS and XP foundation models are projected into a shared embedding space. The projection head for LRS includes a cross-attention block with a learnable query vector, following Parker et al. (2024). For Gaia XP, the projection head is either a cross-attention block (when following the attention-based model) or an MLP (when based on the OAE model). The core alignment objective is the contrastive loss between these projected embeddings.
@@
  • Two cross-modal decoders predict one modality of spectrum from the other.
These components serve to retain modality-specific information, support cross-modal translation, and increase the robustness of the learned representations. Although our multi-decoder design is motivated by the need to enrich spectral embeddings with diverse reconstruction pathways, we note that related ideas have independently appeared in other domains, such as the sensor-agnostic image retrieval framework in remote sensing proposed by Hackstein et al. (2024). As further discussed in the final subsection of this section, reconstructing the original spectra also helps increase the mutual information between the projected embeddings and the inputs.
@@
  • CLIP: A baseline contrastive model with no auxiliary decoders.
@@
2.4 Loss Functions during Contrastive Training
The total training objective $\mathcal{L}_{\text{total}}$ for the SpecCLIP model – whether for the CLIP or CLIP-variants – comprises three components: a contrastive CLIP loss, a reconstruction loss, and a cross-modal prediction loss. The total loss is a weighted sum:
@@ 2.4 Loss Functions during Contrastive Training
where $\delta_{\text{recon}}$ and $\delta_{\text{pred}}$ are binary indicators controlling the inclusion of the reconstruction and cross-modal prediction losses, respectively. The weights $w_{\text{recon}}$ and $w_{\text{pred}}$ control the relative contribution of the reconstruction and prediction losses, and are fixed to 1 (i.e., equal weighting) throughout this work.
@@
The contrastive loss aligns XP and LRS embeddings in a shared embedding space. Let $f_{\text{xp}}$ and $f_{\text{lrs}}$ denote the encoders for XP and LRS inputs (including their pre-trained models and projection networks). For a batch of $N$ paired examples $\{(x_{i}^{\text{xp}},x_{i}^{\text{lrs}})\}_{i=1}^{N}$, their projected embeddings are computed as described below.
In the CLIP-pr model, we use the full projected embeddings and normalize them to unit length: $z_{i}^{\text{xp}}=f_{\text{xp}}(x_{i}^{\text{xp}})$ and $z_{i}^{\text{lrs}}=f_{\text{lrs}}(x_{i}^{\text{lrs}})$, where each embedding is L2-normalized. In the CLIP-split model, only the shared components of the embeddings are used and similarly normalized: $z_{i}^{\text{xp}}=f_{\text{xp}}^{\text{shared}}(x_{i}^{\text{xp}})$ and $z_{i}^{\text{lrs}}=f_{\text{lrs}}^{\text{shared}}(x_{i}^{\text{lrs}})$.
The similarity matrix are scaled by a temperature parameter $\tau$ (set to 15.5, following Parker et al. 2024):
@@
where $\langle\cdot,\cdot\rangle$ denotes the dot product between two embedding vectors. This dot product becomes equivalent to cosine similarity if the vectors are normalized. The parameter $\tau$ controls how confidently the contrastive loss distinguishes positives from negatives. Larger $\tau$ leads to higher confidence and stronger push/pull between matched and mismatched pairs.
@@
where $\ell_{i,:}$ and $\ell_{:,i}$ are the similarity scores for XP-to-LRS and LRS-to-XP matching, respectively, and $\text{CE}(\cdot,i)$ denotes the cross-entropy loss with label $i$ defined as $\text{CE}(\ell_{i,:},i)=-\log\left(\exp(\ell_{i,i})/\sum_{j=1}^{N}\exp(\ell_{i,j})\right)$. The cross-entropy loss encourages matched pairs to have higher similarity than mismatched ones in the shared embedding space.
@@
Each modality has its own decoder to reconstruct the original spectrum from its embedding. In the CLIP-pr model, reconstruction is based on the full projected embedding. In the CLIP-r and CLIP-split model, reconstruction uses both shared and non-shared embeddings, which come from separate projection branches. Let $\hat{x}_{i}^{\text{xp}}$ and $\hat{x}_{i}^{\text{lrs}}$ denote the reconstructions. The loss is:
@@
where $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the LRS input $x_{i}^{\text{lrs}}$, used to normalize the input before reconstruction.
@@
To facilitate spectrum translation, cross-modal decoders predict one modality from the embedding of the other. In CLIP-pr, this is done using the full embedding. In CLIP-p and CLIP-split, only the shared embedding component is used for cross-modal prediction. Let $\hat{x}_{i}^{\text{xp}\leftarrow\text{lrs}}$ and $\hat{x}_{i}^{\text{lrs}\leftarrow\text{xp}}$ be the cross-predicted spectra:
@@
Together, these losses ensure that the model aligns LRS and XP modalities in a shared space (via contrastive loss), retains informative modality-specific content (via reconstruction), and supports spectrum translation across modalities (via prediction). In CLIP-split, this is achieved through a disentangled embedding structure with explicit shared and non-shared components, while CLIP-pr uses a unified embedding vector for all tasks. The loss curves of our CLIP-pr and CLIP-split models can be found in Appendix F.
2.5 Connection of Reconstruction Loss to Mutual Information
Minimizing the reconstruction loss encourages the latent embedding to retain as much information as possible about the original input spectra. This intuition can be formalized via the mutual information between the input spectrum (either $x^{\text{xp}}$ or $x^{\text{lrs}}$) and its corresponding embedding ($z^{\text{xp}}$ or $z^{\text{lrs}}$). For notational simplicity, we use $x$ and $z$ to denote a generic input spectrum and its embedding, respectively. The mutual information can be written as (Barber & Agakov, 2003; Ting, 2025):
@@ 2.5 Connection of Reconstruction Loss to Mutual Information
where the inequality follows from the non-negativity of the Kullback–Leibler divergence $\mathrm{KL}(p(x|z)\|q(x|z))$ (Barber & Agakov, 2003; Poole et al., 2019). Thus, minimizing the expected negative log-likelihood under $q(x|z)$ serves as a variational lower bound on the mutual information $I(z,x)$.
@@ 3 Downstream Tasks
3.1 Models for Parameter Estimation
@@
We explore two complementary approaches for stellar parameter estimation: MLPs and simulation-based inference (SBI), also known as implicit-likelihood or likelihood-free inference (Tejero-Cantero et al., 2020; Ho et al., 2024). SBI enables the inference of parameter posteriors by learning the underlying distribution (e.g., the posterior itself) and evaluating on observed data, without requiring an explicit likelihood function.
3.1.1 Multilayer Perceptrons
@@
For most downstream tasks involving LAMOST LRS and Gaia XP spectra, we employ MLPs due to their training efficiency and scalability, which make them suitable for processing large datasets with limited computational resources. Each MLP has the following layer architecture: $\text{input\_dim},1024,512,64,1]$, where $\text{input\_dim}=\{1462,343,768\}$ depending on whether the input is raw LRS spectra, raw XP spectra, or embeddings. The output dimension is one, corresponding to a single stellar parameter. Each MLP model has approximately 1.4 million trainable parameters when applied to the embeddings.
3.1.2 Simulation-Based Inference
For selected tasks, we also apply SBI for posterior inference. We report the median of the posterior as the point estimate. For pre-trained models via MT, we first reduce the representations by averaging over the sequence dimension (e.g., from $[146,768]$ to $[768]$ for the LAMOST model). The resulting 768-dimensional embeddings and their high-quality labels are then used as input to SBI. For CLIP models and the pre-trained XP model via OAE, no reduction of dimension is required and we directly use the 768-dimensional (projected) embeddings together with their labels.
@@ 3.1.2 Simulation-Based Inference
We follow the Neural Posterior Estimation (NPE) framework to directly estimate the posterior distribution of parameters given observations, using neural density estimators. This NPE-based approach is relatively straightforward to implement and computationally efficient, making it a practical choice compared to other variants such as Neural Likelihood Estimation (NLE). We adopt two types of (conditional) normalizing flows as density estimators from the sbi package (Tejero-Cantero et al., 2020): Masked Autoregressive Flow (MAF) (Papamakarios et al., 2017) and Neural Spline Flow (NSF) (Durkan et al., 2019). Each SBI model uses two transformations with 60 hidden units (unless otherwise noted), totaling roughly 0.1 million trainable parameters. Details of the adopted normalizing flows are descrbed in Appendix C.
@@
To evaluate the calibration of the inferred posteriors, we perform simulation-based calibration (SBC, Talts et al., 2018) using rank statistics, also known as Probability Integral Transform (PIT) values in some literature (Gneiting et al., 2007), complemented by the Kolmogorov–Smirnov (K–S) test. For each (out of 200 in total) spectrum–parameter pair, we draw posterior samples and compute the rank of the ground-truth parameter value. Well-calibrated posteriors yield uniformly distributed ranks, which we assess via the K–S test. We consider results valid if the $p$-value exceeds 0.05, indicating no significant deviation from uniformity and hence reliable posterior coverage.
3.2 Sample Selection for Parameter Estimation
@@
We select approximately 100,000 stars per parameter to balance the parameter distribution and computational tractability. The quality-control criteria include $g$-band spectral SNRg >> 20, Gaia $g$-band magnitude << 16.5, and $|v_{r}|<800$ km s-1 to exclude likely extragalactic sources. LAMOST DR11 is used throughout; we adopt a $3^{\prime\prime}$ matching radius between it and other catalogs. For APOGEE, we typically (unless otherwise noted) require APOGEE SNR (median SNR per pixel in combined frame (at apStar sampling))>> 50.
@@
  • Radial Velocity ($v_{r}$): We select 100,299 stars in common between APOGEE DR17 (Abdurro’uf et al., 2022) and the LAMOST LRS sample. Their radial velocities are approximately uniformly distributed over the range [800,800][-800,800] km s-1, divided into 800 bins, each containing up to 1600 stars.
@@
  • Effective Temperature ($T_{\rm eff}$): Approximately 100,000 stars are uniformly sampled in the $T_{\rm eff}$$\log g$ plane, divided into a $100\times 100$ grid, with up to four stars per bin. Values of $T_{\rm eff}$ are adopted from the LAMOST catalog.
@@
  • Surface Gravity ($\log g$): 100,000 stars combining log $g$ labels from Kepler (Li et al., 2022b, mainly for red giant stars) and APOGEE DR17 (Abdurro’uf et al., 2022). Binned into 750 intervals (up to 350 stars per bin), with priority given to Kepler-based values.
@@
  • Iron Abundance ($\mathrm{[Fe/H]}$): A total of 100,118 stars are selected from a merged dataset comprising APOGEE DR17 (Abdurro’uf et al., 2022, for stars with ${\rm[Fe/H]}>-2.0$), supplemented with metal-poor stars from the PASTEL and SAGA compilations (Huang et al., 2024), the LAMOST/Subaru VMP sample (Li et al., 2022a), and UMP datasets (Sestito et al., 2019). A targeted sampling strategy is employed to ensure balanced coverage in [Fe/H].
@@
  • Other Elemental Abundances: 85,400 stars from LAMOST–APOGEE cross-matches, filtered by APOGEE SNR >> 40 and valid abundance flags (i.e., C_FE_FLAG = 0, N_FE_FLAG = 0, MG_FE_FLAG = 0). A 2D binning in [Mg/Fe]–[Fe/H] space ($632\times 632$ bins) ensures broad and uniform sampling. The $\alpha$/Fe] values used here are computed from APOGEE measurements as ALPHA_M - (FE_H - M_H), where ALPHA_M includes O, Mg, Si, S, Ca, Ti, and Ti II (Jönsson et al., 2020).
@@
  • Extinction (E(BPRP)E(BP-RP)): A total of 86,000 stars, with extinction values estimated using the star-pair technique (Yuan et al., 2013) based on the LAMOST stellar parameter catalog.
@@
  • Asteroseismic Parameters: A total of 3,029 stars have asteroseismic parameters $\Delta\nu$ and $\nu_{\rm max}$ derived from Kepler light curves (Li et al., 2022b; Chaplin et al., 2014), including 2,718 red giant stars and 311 main-sequence/turn-off stars. In addition, 4,034 red giant stars have measured gravity-mode period spacing $\Delta\Pi$, taken from published catalogs (Vrard et al., 2016).
@@
For Gaia XP spectra, we examine the following key parameters (with sample number in brackets) – $\alpha$/Fe] (94,584), [C/Fe] (99,934), [N/Fe] (95,089), $T_{\rm eff}$ (100,000), log $g$, [Fe/H] (113,218), and color excess E(BPRP)E(BP-RP) (99,087) – based on consistent data sources. These datasets are cross-matched with the Gaia XP catalog. To enhance [C/Fe] coverage in the metal-poor regime, we further include stars from the LAMOST very metal-poor catalog, where [C/Fe] has been estimated using a customized version of the SEGUE Stellar Parameter Pipeline (LSSPP; Lee et al., 2015).
@@
All datasets are split in a ratio of 0.81:0.09:0.10 for training, validation, and testing, respectively. The validation set is used for early stopping, that is, training is halted when performance on the validation set no longer improves within 10 training epochs. The testing set is held out and used exclusively for reporting all downstream task results shown in the tables throughout the paper. An exception is made for the figures generated from MLP-based downstream models, where we combine the training and validation sets for model training, as explained in Section 5.5.

### v4

@@ II.2 Pre-trained Foundation Model for Gaia XP Spectra
Gaia XP low-resolution spectra (2023A&A...674A...2D), obtained via the Blue (with resolving power $R$ ranging from 30 to 100) and Red ($R$ ranging from 70 to 100) Photometers (BP and RP), are essential for determining key stellar properties such as stellar atmospheric parameters and chemical compositions. The differences in resolution and wavelength range compared to LAMOST LRS motivate the construction of a separate foundation model tailored to the XP modality.
@@
Two types of models are explored for Gaia XP. The first is a transformer-based MT model, structurally similar to that used for LAMOST LRS, but with tokenization at the individual flux-point level, yielding 343 tokens per spectrum, together with two special tokens—the mean and standard deviation of the spectrum. Like the LRS model, this version uses 6 self-attention layers and masks 6 chunks (20 tokens each, corresponding to an overall masking rate of $\approx$ 35%) during training.
@@
The second model is a MLP-based ordinary autoencoder (denoted as OAE), with a bottleneck layer of 768 dimensions. Both XP models have approximately the same number of trainable parameters as the LRS model and are trained for the same number of epochs. In this paper, we use the OAE for Gaia XP to test results in tables, unless otherwise noted, because of its better performance, as discussed in Section VI.6.
II.3 Contrastive Learning with Decoders
@@
To align the two spectral modalities, we use 820,568 paired LAMOST LRS and Gaia XP spectra. The backbone of the alignment model leverages contrastive loss, augmented by auxiliary decoders that contribute to additional supervision.
@@
As shown in Figure 1, the embeddings of the LRS and XP foundation models are projected onto a shared embedding space. The LRS projection head includes a cross-attention block with a recurrence vector that is learnable, following 2024MNRAS.531.4990P. For Gaia XP, the projection head is either a cross-attention block (when following the attention-based model) or an MLP (when based on the OAE model). The core alignment objective is the contrastive loss between these projected embeddings.
@@
  • Two cross-modal decoders predict one spectrum modality from the other.
These components serve to retain modality-specific information, support cross-modal translation, and increase the robustness of learned representations. Although our multi-decoder design is motivated by the need to enrich spectral embeddings with diverse reconstruction pathways, we note that related ideas have appeared independently in other domains, such as the sensor-agnostic image retrieval framework in remote sensing proposed by 2024arXiv240107782H. As discussed in further detail in the final subsection of this section, reconstructing the original spectra also helps to increase the mutual information between the projected embeddings and the inputs.
@@
  • CLIP: A baseline contrastive model without auxiliary decoders.
@@
II.4 Loss Functions during Contrastive Training
The total training objective $\mathcal{L}_{\text{total}}$ for the SpecCLIP model – whether for CLIP or CLIP-variants – comprises three components: a contrastive CLIP loss, a reconstruction loss, and a cross-modal prediction loss. The total loss is a weighted sum:
@@ II.4 Loss Functions during Contrastive Training
where $\delta_{\text{recon}}$ and $\delta_{\text{pred}}$ are binary indicators that control the inclusion of reconstruction losses and cross-modal prediction, respectively. The weights $w_{\text{recon}}$ and $w_{\text{pred}}$ control the relative contribution of the reconstruction and prediction losses and are fixed at 1 (i.e., equal weighting) throughout this work.
@@
The contrastive loss aligns the XP and LRS embeddings in a shared embedding space. Let $f_{\text{xp}}$ and $f_{\text{lrs}}$ denote the encoders for XP and LRS inputs (including their pre-trained models and projection networks). For a batch of $N$ paired examples $\{(x_{i}^{\text{xp}},x_{i}^{\text{lrs}})\}_{i=1}^{N}$, their projected embeddings are computed as described below.
In the CLIP-pr model, we use all projected embeddings and normalize them to unit length: $z_{i}^{\text{xp}}=f_{\text{xp}}(x_{i}^{\text{xp}})$ and $z_{i}^{\text{lrs}}=f_{\text{lrs}}(x_{i}^{\text{lrs}})$, where each embedding is L2-normalized. In the CLIP-split model, only the shared components of the embeddings are used and similarly normalized: $z_{i}^{\text{xp}}=f_{\text{xp}}^{\text{shared}}(x_{i}^{\text{xp}})$ and $z_{i}^{\text{lrs}}=f_{\text{lrs}}^{\text{shared}}(x_{i}^{\text{lrs}})$.
The similarity matrix are scaled by a temperature parameter $\tau$ (set to 15.5, following 2024MNRAS.531.4990P):
@@
where $\langle\cdot,\cdot\rangle$ denotes the dot product between two embedding vectors. This dot product becomes equivalent to cosine similarity if the vectors are normalized. The parameter $\tau$ controls how confidently the contrastive loss distinguishes positives from negatives. The larger $\tau$ leads to greater confidence and greater push/pull between the matched and mismatched pairs.
@@
where $\ell_{i,:}$ and $\ell_{:,i}$ are the similarity scores for XP-to-LRS and LRS-to-XP matching, respectively, and $\text{CE}(\cdot,i)$ denotes the cross-entropy loss with label $i$ defined as $\text{CE}(\ell_{i,:},i)=-\log\left(\exp(\ell_{i,i})/\sum_{j=1}^{N}\exp(\ell_{i,j})\right)$. The cross-entropy loss encourages the matched pairs to have a higher similarity than the mismatched ones in the shared embedding space.
@@
Each modality has its own decoder to reconstruct the original spectrum from its embedding. In the CLIP-pr model, reconstruction is based on the full projected embedding. In the CLIP-r and CLIP-split models, reconstruction uses both shared and non-shared embeddings, which come from separate projection branches. Let $\hat{x}_{i}^{\text{xp}}$ and $\hat{x}_{i}^{\text{lrs}}$ denote the reconstructions. The loss is:
@@
where $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the LRS input $x_{i}^{\text{lrs}}$, used to further normalize the input before reconstruction [^4] This normalization step is optional. We initially adopted this design for potential training stability. Although each spectrum is already flux-normalized by continuum division during pre-processing, this additional normalization (1) centers the input distribution and (2) focuses the decoder on learning the shape of spectral features. This makes the projected embedding retain essential LRS information. In contrast, for Gaia XP, we reconstruct the relative fluxes (colors) normalized by the 550 nm flux, aiming to recover the full chromatic information..
@@
To facilitate spectrum translation, cross-modal decoders predict one modality from the embedding of the other. In CLIP-pr, this is done by using the full embedding. In CLIP-p and CLIP-split, only the shared embedding component is used for cross-modal prediction. Let $\hat{x}_{i}^{\text{xp}\leftarrow\text{lrs}}$ and $\hat{x}_{i}^{\text{lrs}\leftarrow\text{xp}}$ be the cross-predicted spectra:
@@
II.5 Connection of Reconstruction Loss to Mutual Information
Minimizing the reconstruction loss encourages the latent embedding to retain as much information as possible about the original input spectra. This intuition can be formalized via the mutual information between the input spectrum (either $x^{\text{xp}}$ or $x^{\text{lrs}}$) and its corresponding embedding ($z^{\text{xp}}$ or $z^{\text{lrs}}$). For notational simplicity, we use $x$ and $z$ to denote a generic input spectrum and its embedding, respectively. The mutual information can be written as (10.5555/2981345.2981371; 2025arXiv250612230T):
@@ II.5 Connection of Reconstruction Loss to Mutual Information
where the inequality follows from the non-negativity of the Kullback–Leibler divergence $\mathrm{KL}(p(x|z)\|q(x|z))$ (10.5555/2981345.2981371; 2019arXiv190506922P). Thus, minimizing the expected negative log-likelihood under $q(x|z)$ serves as a variational lower bound on the mutual information $I(z,x)$.
@@ III Downstream Tasks
III.1 Models for Parameter Estimation
@@
We explore two complementary approaches for stellar parameter estimation: MLPs and simulation-based inference (SBI), also known as implicit-likelihood or likelihood-free inference (tejero-cantero2020sbi; 2024OJAp....7E..54H). SBI enables inference of parameter posteriors by learning the underlying distribution (e.g., the posterior itself) and evaluating on observed data, without requiring an explicit likelihood function.
III.1.1 Multilayer Perceptrons
@@
For most downstream tasks involving LAMOST LRS and Gaia XP spectra, we employ MLPs due to their training efficiency and scalability, which make them suitable for processing large datasets with limited computational resources. Each MLP has the following layer architecture: $\text{input\_dim},1024,512,64,1]$, where $\text{input\_dim}=\{1462,343,768\}$ depending on whether the input is raw LRS spectra, raw XP spectra, or embeddings. The output dimension is one, corresponding to a single stellar parameter. For pre-trained models via MT, we first reduce the representations by averaging over the sequence dimension (e.g., from $[146,768]$ to $[768]$ for the LAMOST model). The resulting 768-dimensional embeddings and their high-quality labels are then used as input to the MLP. For CLIP models and the pre-trained XP model via OAE, no reduction of dimension is required and we directly use the 768-dimensional (projected) embeddings together with their labels. Each MLP model has approximately 1.4 million trainable parameters when applied to the embeddings.
III.1.2 Simulation-Based Inference
For selected tasks, we also apply SBI for posterior inference. We report the median of the posterior as the point estimate.
@@ III.1.2 Simulation-Based Inference
We follow the Neural Posterior Estimation (NPE) framework to directly estimate the posterior distribution of parameters given observations, using neural density estimators. This NPE-based approach is relatively straightforward to implement and computationally efficient, making it a practical choice compared to other variants such as Neural Likelihood Estimation (NLE). We adopt two types of (conditional) normalizing flows as density estimators from the sbi package (tejero-cantero2020sbi): Masked Autoregressive Flow (MAF) (2017arXiv170507057P) and Neural Spline Flow (NSF) (2019arXiv190604032D). Each SBI model uses two transformations with 60 hidden units (unless otherwise noted), totaling roughly 0.1 million trainable parameters. Details of the adopted normalizing flows are described in Appendix C.
@@
To evaluate the calibration of the inferred posteriors, we perform simulation-based calibration (SBC, talts2018validating) using rank statistics, also known as Probability Integral Transform (PIT) values in some literature (PITvalue), complemented by the Kolmogorov–Smirnov (K–S) test (Kolmogorov1992). The K–S test measures the maximum discrepancy between the empirical cumulative distribution function (CDF) of the ranks and the expected uniform CDF, providing a non-parametric test of distributional consistency. For each (out of 200 in total) spectrum–parameter pair, we draw posterior samples and compute the rank of the ground-truth parameter value. Well-calibrated posteriors yield uniformly distributed ranks, which we assess via the K–S test. We consider results valid if the $p$-value exceeds 0.05, indicating that there is no significant deviation from uniformity and therefore reliable posterior coverage.
IV Datasets
IV.1 Sample Selection for Pre-training
For the LRS model, we select a subset of 966,082 high-quality spectra for pre-training, using a 9:1 train–validation split (the same split ratio is used for the other modeling efforts described in the following two subsections). The selection criteria are: (1) SNR in the SDSS $g$-band greater than 50, and (2) apparent $g$-band magnitude less than 15.8. Additionally, we include all spectral types beyond the AFGK classes, while attempting to balance the four AFGK types themselves. We also aim to balance giant and dwarf stars; however, due to the intrinsic distribution of the dataset, the final ratio of dwarfs to giants is approximately 3.8:1.
For the XP model, we pre-train this model using one million Gaia XP spectra, with around 80% having matching LAMOST LRS spectra. Each XP spectrum consists of 343 flux points spanning wavelengths from 336 nm to 1021 nm.
IV.2 Sample Selection for Parameter Estimation
@@
We selected approximately 100,000 stars per parameter to balance the parameter distribution and computational tractability. The quality-control criteria include $g$-band spectral SNRg >> 20, Gaia $g$-band magnitude << 16.5, and $|v_{r}|<800$ km s-1 to exclude likely extragalactic sources. LAMOST DR11 is used throughout; we adopt a $3^{\prime\prime}$ matching radius between it and other catalogs. For APOGEE, we typically (unless otherwise noted) require APOGEE SNR (median SNR per pixel in combined frame (at apStar sampling))>> 50.
@@
  • Radial Velocity ($v_{r}$): We select 100,299 stars in common between APOGEE DR17 (2022ApJS..259...35A) and the LAMOST LRS sample. Their radial velocities are approximately uniformly distributed over the range [800,800][-800,800] km s-1, divided into 800 bins, each containing up to 1600 stars.
@@
  • Effective Temperature ($T_{\rm eff}$): Approximately 100,000 stars are uniformly sampled in the $T_{\rm eff}$$\log g$ plane, divided into a $100\times 100$ grid, with up to four stars per bin. The values of $T_{\rm eff}$ are adopted from the LAMOST catalog.
@@
  • Surface Gravity ($\log g$): 100,000 stars combining log $g$ labels from Kepler (2022ApJ...927..167L, mainly for red giant stars) and APOGEE DR17 (2022ApJS..259...35A). Binned into 750 intervals (up to 350 stars per bin), with priority given to Kepler-based values.
@@
  • Iron Abundance ($\mathrm{[Fe/H]}$): A total of 100,118 stars are selected from a merged dataset comprising APOGEE DR17 (2022ApJS..259...35A, for stars with ${\rm[Fe/H]}>-2.0$), supplemented with metal-poor stars from the PASTEL and SAGA compilations (2024ApJ...974..192H), the LAMOST/Subaru VMP sample (2022ApJ...931..147L), and UMP datasets (2019MNRAS.484.2166S). A targeted sampling strategy is used to ensure balanced coverage in [Fe/H].
@@
  • Other Elemental Abundances: 85,400 stars from LAMOST–APOGEE cross-matches, filtered by APOGEE SNR >> 40 and valid abundance flags (i.e., C_FE_FLAG = 0, N_FE_FLAG = 0, MG_FE_FLAG = 0). A 2D binning in [Mg/Fe]–[Fe/H] space ($632\times 632$ bins) ensures broad and uniform sampling. The $\alpha$/Fe] values used here are computed from APOGEE measurements as ALPHA_M - (FE_H - M_H), where ALPHA_M includes O, Mg, Si, S, Ca, Ti, and Ti II (2020AJ....160..120J).
@@
  • Extinction (E(BPRP)E(BP-RP)): A total of 86,000 stars, with extinction values estimated using the star-pair technique (2013MNRAS.430.2188Y) based on the LAMOST stellar parameter catalog.
@@
  • Asteroseismic Parameters: A total of 3,029 stars have asteroseismic parameters $\Delta\nu$ and $\nu_{\rm max}$ derived from Kepler light curves (2022ApJ...927..167L; 2014ApJS..210....1C), including 2,718 red giant stars and 311 main-sequence/turn-off stars. In addition, 4,034 red giant stars have measured gravity-mode period spacing $\Delta\Pi$, taken from published catalogs (2016A&A...588A..87V).
@@
For Gaia XP spectra, we examine the following key parameters (with sample number in brackets) – $\alpha$/Fe] (94,584), [C/Fe] (99,934), [N/Fe] (95,089), $T_{\rm eff}$ (100,000), log $g$, [Fe/H] (113,218), and color excess E(BPRP)E(BP-RP) (99,087) – based on consistent data sources. These datasets are cross-matched with the Gaia XP catalog. To enhance [C/Fe] coverage in the metal-poor regime, we further include stars from the LAMOST very metal-poor catalog, where [C/Fe] has been estimated using a customized version of the SEGUE Stellar Parameter Pipeline (LSSPP; 2015AJ....150..187L).
@@
All datasets are divided into a ratio of 0.81:0.09:0.10 for training, validation, and testing, respectively. The validation set is used for early stopping, that is, training is halted when performance on the validation set no longer improves within 10 training epochs. The testing set is kept out and used exclusively for reporting all downstream task results shown in the tables throughout the paper. An exception is made for the figures generated from the MLP-based downstream models, where we combine the training and validation sets for model training, as explained in Section VI.5.
IV.3 Sample Selection for External Validation
We use multiple external datasets for validation. For LAMOST LRS, we perform two types of comparisons:
  1. 1.
    Radial Velocity: Compared with GALAH DR4 (2025PASA...42...51B), where 49,905 stars are selected for which GALAH’s global RV fit succeeded (FIT_GLOBAL_RV = True), with SNR_PX_CCD2 $\geq$ 20, LAMOST SNRg $\geq$ 20, and both GALAH RV_COMP_1 and LAMOST RV having absolute values $\leq$ 999 km s-1.
  2. 2.
    Iron Abundance:
    • DESI DR1 (2025arXiv250314745D; 2025arXiv250514787K): 119,335 stars are selected with LAMOST SNRg $\geq$ 30, DESI SN_B $\geq$ 30, SUCCESS = True, and [Fe/H] $\geq-3.8$.
    • GALAH DR4: 33,411 stars are retained with FLAG_FE_H = 0, SNR_PX_CCD2 $\geq$ 30, LAMOST SNRg $\geq$ 30 in both surveys, and $T_{\rm eff}$ $\geq$ 4000 K.
For Gaia XP, the iron abundance validation uses:
  1. 1.
    GALAH DR4: 33,411 stars satisfying FLAG_SP = 0, SNR_PX_CCD2 $\geq$ 50, and $T_{\rm eff}$ $\geq$ 4000 K.
  2. 2.
    Gaia RVS catalog (2024A&A...683L..11V): 1,413 cross-matched stars are used.
IV.4 Preprocessing
For the LAMOST LRS spectrum, to focus on the most informative spectral features, we retain the 400–560 nm wavelength range, resulting in 1462 flux points per spectrum. The spectra are normalized before being input into the model; the normalization procedure is described in Appendix B, where an iterative polynomial fitting algorithm robustly estimates the stellar continuum across both blue and red wavelength segments while suppressing absorption features and noise. For the Gaia XP spectra, broad-band color information may sometimes carry more discriminative power than individual spectral features. We therefore normalize each spectrum by its flux at 550 nm which lies near the center of the $V$-band.